OpenStudy (anonymous):

is 3(12x+2)=36x-9 an infinitely many number?

OpenStudy (jdoe0001):

"an infinitely many number?" looks very limited in characters there... so I'd think not, since it only has 15 characters

OpenStudy (anonymous):

3(12x + 2) = 36x - 9 --- distribute through the parenthesis 36x + 6 = 36x - 9 --- subtract 36x 36x - 36x + 6 = -9 -- now subtract 6 36x - 36x = -9 - 6 -- combine like terms 0 = - 15 (incorrect) There is NO SOLUTION to this problem

OpenStudy (anonymous):

thanks you both.